Grade 12Mechanics

What is the minimum velocity required for a projectile to pass through a point (a,b) starting from origin.no angle is given.

Srilekha
8 Years agoGrade 12

1 Answer

Khimraj
8 Years ago
there will be minimum velocity required if maximum height point is (a,b)then sin2\Theta = b2/(a2+ b2)and max height is given byh = vo2sin2\Theta/2g2bg =vo2b2/(a2+ b2)So vo2 = 2g(a2+b2)/band vo = \sqrt{2g(a^{2}+b^{2})/b}